OpenStudy (bduke4):

How many grams of Cl are in 275 g of CaCl2

OpenStudy (jazmizzlemynizzle):

Okay so if CF2Cl2 has a mass of 55g and molar mass of 12.01+(2*18.00)+(2*35.45)= 118.91g/mol Using the formula n of CF2Cl2= mass of CF2Cl2/molar mass of CF2Cl2 55g/118.91g/mol = 0.46 mol So using this you can see that 0.46 mol of CF2Cl2 must be present using this mol we can find CL2 in moles n=mass of Cl/molar mass of total Cl rearrange it molar mass of total Cl2*n = mass of Cl So in conclusion - 0.46mol * (2*35.45g/mol)=33g of Cl in CF2Cl2